OpenStudy (anonymous):

Suppose you want to show the data in a circle graph. Find the measure of the central angle for 7th grade. Round to the nearest degree.

OpenStudy (anonymous):

This is the table for the problem.

OpenStudy (anonymous):

63.144 degrees. So the answer is 63.

OpenStudy (anonymous):

Thanks. How did you get it?

OpenStudy (anonymous):

Do you need to show work?

OpenStudy (anonymous):

No. I was just wondering for my own knowledge. I wanted to know the process so I can understand the problem.

OpenStudy (anonymous):

Find the percentage of 7th graders. (50/285)*100 = 17.54%

OpenStudy (anonymous):

Take 17.54% of 360 17.54% of 360 = 0.1754*360 = 63.144

OpenStudy (anonymous):

Thanks. I really appreciate it. c:

OpenStudy (anonymous):

No problem mate. Glad to help.